OpenStudy (anonymous):

What is the exact area of a circle with diameter 1 mile? pi divided by 4 square miles pi divided by 2 square miles pi square miles 2 times pi square miles

OpenStudy (mertsj):

If the diameter is 1, what is the radius?

OpenStudy (anonymous):

a = pi * r^2 d = 2r 1 mi = 2r r = .5 mi a = pi * .25 mi

OpenStudy (anonymous):

i dont know?

OpenStudy (anonymous):

can you show me how to do it

OpenStudy (mertsj):

Divide the diameter by 2 to find the radius.

OpenStudy (mertsj):

1 divided by 2 is 1/2 and that is the radius. Now square the radius. What is (1/2)^2 ???

OpenStudy (anonymous):

0.25

OpenStudy (mertsj):

yes. 1/4 Now to find the area multiply 1/4 by pi so : \[A=\frac{\pi}{4}\]

OpenStudy (mertsj):

Now pick the right answer from your list.

OpenStudy (anonymous):

ok thanks

OpenStudy (mertsj):

What did you pick?

OpenStudy (anonymous):

its π4

OpenStudy (mertsj):

1, 2, 3, or 4 ???

OpenStudy (anonymous):

it's pi / 4 mi^2

OpenStudy (anonymous):

it's an alias for pi * .25 mi^2, since .25 = 1/4, pi * 1/4 = pi / 4

OpenStudy (anonymous):

its 4

OpenStudy (mertsj):

"divided by" means "over"

OpenStudy (mertsj):

So now which is the correct answer

OpenStudy (anonymous):

pi divided by 4 square miles

OpenStudy (mertsj):

That's better.

OpenStudy (anonymous):

thanks

OpenStudy (mertsj):

yw